OpenStudy (anonymous):

Need to check an integral answer. integral sin^5(x) * cos^4(x) dx

OpenStudy (anonymous):

I said u = cos(x) so du/-sin(x) = dx

OpenStudy (anonymous):

so now I have...

OpenStudy (anonymous):

\[\int\limits (1-\cos^2(x))(1-\cos^2(x))(\cos^4(x))(\sin(x)\] dx

OpenStudy (anonymous):

after substitution I have ...

OpenStudy (anonymous):

\[\int\limits (1-u^2)(1-u^2)(u^4)(\sin(x))du/-\sin(x)\]

OpenStudy (anonymous):

which should simplify to...

OpenStudy (anonymous):

\[\int\limits u^8 - 2u^6 + u^4 du\]

OpenStudy (anonymous):

final answer (-1/9)cos^9(x) + (2/7)cos^7(x) + (1/5)cos^5(x) + C

OpenStudy (anonymous):

What do you think John?

OpenStudy (johnweldon1993):

Doing it out on paper, hang on 1 sec

OpenStudy (anonymous):

Thank you.

OpenStudy (johnweldon1993):

Ehh by parts is terrible...let me try a reduction instead sorry for the wait @jasonjohnson86

OpenStudy (anonymous):

It should just be either a sin or cos sub (; Thats what Im working out of in my book.

OpenStudy (anonymous):

With the power reduction.

OpenStudy (johnweldon1993):

Yeah I thought somewhere along the way something got left behind...just trying to figure out where... And I just check wolfram just to double check http://www.wolframalpha.com/input/?i=integral+%28sin%5E5%28x%29cos%5E4%28x%29dx%29

OpenStudy (anonymous):

Yeah already checked that, don''t like the form of it. I think I should be able to get an answer this route.

OpenStudy (anonymous):

Yeah this should be good.....unless anyone has any issues with what I did. I dropped a negative sign on one term. should be (-1/5)cos^5(x)

OpenStudy (anonymous):

solution should be (-1/9)cos^9(x) + (2/7)cos^7(x) + (-1/5)cos^5(x)

OpenStudy (primeralph):